1.two cars are 500 miles apart and moving directly towards each other.One car is moving at a speed of 100 mph and the other is moving at 70 mph.Assuming that the cars start moving at the same time,how long does it take for the two cars to meet? Repeat the previous example except this time assume that the faster car will start 1 hour after slower car starts?
2.two boats start out 100 miles apart and start moving to the right at the same time.the boat on the left is moving at twice the speed as the boat on the righ.Five hours after starting,the boat on the left catches up with the boat on the right.How fast was each boat?

Answer by stanbon(75887)

1.two cars are 500 miles apart and moving directly towards each other.One car is moving at a speed of 100 mph and the other is moving at 70 mph.Assuming that the cars start moving at the same time,how long does it take for the two cars to meet?
----
The closing speed of the two cars is (100 + 70)mph
Equation:
Time = distance/rate = 500/170 = 2.94 hrs
---------------------------------------------
Repeat the previous example except this time assume that the faster car will start 1 hour after slower car starts?
Slower car has gone 70 miles in the 1st hour leaving 430 miles to be covered by
both.
Time = 430/170 = 2.53 hrs.
=========================================


2.Two boats start out 100 miles apart and start moving to the right at the same time.the boat on the left is moving at twice the speed as the boat on the right.
Five hours after starting,the boat on the left catches up with the boat on the right.
How fast was each boat?
---
Left boat DATA:
rate = 2x mph ; time = 5 hrs. ; distance = rt = 10x miles
---
Right boat DATA:
rate = x mph ; time = 5 hrs. ; distance = 5x
-----
Equation:
right boat distance + 100 mi = left boat distance
5x + 100 = 10x
5x = 100
x = 20 mph (right boat speed)
2x = 40 mph (left boat speed)
==================================
Cheers,
Stan H.